A 14 kg hanging sculpture is suspended by a 95-cm-long, 6.0 g steel wire. When the wind blows hard, the wire hums at its fundamental frequency. What is the frequency of the hum?
Answer #1

The warelength of the ware in the string. X=2L=2(0.95) = 19m T= Mg = 14(48) = 137-2N kg/m - m - 0.006 0.95 6.31x103 H 11 137.